ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

5. Differential Geometry and Global Analysis

1991 MS Classification: 53C Hideya Hashimoto, Nippon Institute of technology Grassmann Geometry of 6-dimensional sphere In this talk, we shall give the examples of submanifolds which are invariant under the G_{2} invariant almost Hermitian structure of S^{6}. First we shall give the classication of the Grassmann subbundles of such submanifolds. Next we shall discuss the existence and obstructions of such submanifolds of S^{6}.